# A Simple Baseline for Supervised Surround-view Depth Estimation

### v3

@@ III-F Loss Function
$L_{depth}=\frac{1}{N_{valid}}\sum_{i}\mathbb{(}{D}^{gt}>0)\left|{D}^{pred}_{i}-{D}^{gt}_{i}\right|$ (4)
where $N_{valid}$ stands for the number of valid pixels, $({D}^{gt}>0)$ indicates whether a pixel has the corresponding point cloud for ground-truth depth, ${D}^{pred}$ denotes the predicted depth, and ${D}^{gt}$ represents the ground-truth depth.
@@
$L_{smooth}=\sum_{x}\left|\partial_{x}d^{\prime}_{x}\right|e^{-\left\|\partial_{x}I_{x}\right\|}+\left|\partial_{y}d^{\prime}_{x}\right|e^{-\left\|\partial_{y}I_{x}\right\|}$ (5)
@@
$L=L_{depth}+\lambda L_{smooth}$ (6)
@@ IV Experiments
IV-A Datasets
We conduct experiments on both the nuScenes datasets [34] and Dense Depth for Automated Driving (DDAD) [33].
NuScenes [34]. The nuScenes benchmark is an up-to-date dataset for 3D object detection, as well as semantic segmentation, which is captured in both Boston and Singapore by using a capture system with six cameras, one LiDAR, and five radars. It collects 1000 scenes which are divided into 700/150/150 scenes for training/validation/testing. The ground truth depths are generated by processing the LiDAR point cloud. However, conducting supervised depth on the nuScenes dataset is difficult because of the sparse point cloud and relatively small overlap across cameras. Thus, it has not been applied as the benchmark for supervised surround-view depth evaluation so far. We introduce a simple baseline, which is robust enough to achieve state-of-the-art performance on this dataset by a large margin, compared with previous methods. During the training and evaluation procedure, we consider the max distance $D_{max}$ as 80m.
DDAD [33]. The DDAD dataset is a large-scale dataset for long-range (up to 250m) and dense depth estimation in challenging conditions. The dataset is collected by six cameras and a high-resolution Luminar-H2, which is used to generate dense ground-truth depth maps. There are 12,650 training samples (75,900 images) and 3,950 validation samples (15,800 images) in this dataset. Compared to other related methods, S3Depth can first establish a baseline for surround-view depth estimation. During the training and evaluation procedure, we consider the max distance $D_{max}$ as 200m.
Evaluation Metrics. We assess the results of our method following the evaluation protocol of the depth estimation task [60, 33, 40, 41]. Specifically, we report seven standard metrics (mean Absolute Relative error (Abs Rel), mean Squared Relative error (Sq Rel), Root Mean Squared Error (RMSE), RMSE log, $\delta<1.25$, $\delta<1.25^{2}$ $\delta<1.25^{3}$). While evaluating self-supervised depth estimation methods, the scores are averaged per-frame with median-scaling [40, 41].
@@ IV-B Implementation Details
BTS We use the official implementation of BTS [76]. Densenet161 [77] is used as the backbone. The number of epochs is set to 10 for nuScenes and DDAD. Silog is employed as the loss function.
PixelFormer We use the official implementation of PixelFormer [35]. Swin-Tiny [66] is used as the backbone. The number of epochs is set to 10 for nuScenes and 20 for DDAD. Silog is employed as the loss function.
SurroundDepth [41] We use the official implementation of SurroundDepth [41], and only change the supervision from self-supervised to LiDAR supervised. L1 loss is employed for training the model.
S3Depth We conduct experiments on both DDAD [33] and nuScenes [34] datasets for depth estimation. S3Depth is implemented with the deep-learning library PyTorch [78]. We employ MPViT-small with ImageNet [79] pre-trained weight as the encoder. At each scale, we use $Z=8$ adjacent-view attention layers. Eight NVIDIA 3090 24GB GPUs are used to train our model. The value of $\lambda$ is set to 0.01 in this paper. Adam [80] is employed as our optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.999$ on both datasets. For the nuScenes [34] datasets, we train the S3Depth with batch size 48, learning rate 1e-4 for 10 epochs. The resolution of input images is downsampled to 640$\times$352. When evaluating the models, the resolution of the predicted depth map is upsampled to 900$\times$1600. For the DDAD dataset which is larger than nuScenes, we train the S3Depth with batch size 48, learning rate 1e-4 for 20 epochs. The resolution of input images is downsampled to $640\times 384$ on DDAD. During the evaluation, the resolution of the predicted depth map is upsampled to $1216\times 1936$.
TABLE I: Quantitative depth estimation results on the nuScenes[34] val set. The symbol $*{}^{*}$ indicates the implementation of SurroundDepth [41]. Resolution refers to the resolution of evaluation. The symbol ${}^{\dagger}$ indicates our implementations. S denotes self-supervised, D denotes depth supervision with Lidar and MS indicates that the scores are averaged per-frame with median-scaling [40, 41] at test time.
Method Resolution Train Test lower is better higher is better
@@
$*{}^{*}$ [60] 900 $\times$ 1600 S MS 0.287 3.349 7.184 0.345 0.641 0.845 0.925
$*{}^{*}$ [33] 900 $\times$ 1600 S MS 0.309 2.891 7.994 0.390 0.547 0.796 0.899
FSM [40] 900 $\times$ 1600 S MS 0.299 - - - - - -
$*{}^{*}$ [40] 900 $\times$ 1600 S MS 0.334 2.845 7.786 0.406 0.508 0.761 0.894
$*{}^{*}$ [41] 900 $\times$ 1600 S MS 0.245 3.067 6.835 0.321 0.719 0.878 0.935
${}^{\dagger}$ [76] 900 $\times$ 1600 D - 0.128 0.788 3.750 0.199 0.862 0.943 0.971
${}^{\dagger}$ [35] 900 $\times$ 1600 D - 0.099 0.553 3.171 0.165 0.904 0.959 0.979
${}^{\dagger}$ [41] 900 $\times$ 1600 D - 0.158 1.189 5.028 0.239 0.813 0.930 0.965
Ours 900 $\times$ 1600 D - 0.067 0.673 2.457 0.144 0.951 0.970 0.981
TABLE I: Quantitative depth estimation results on the nuScenes[34] val set. The symbol $*{}^{*}$ indicates the implementation of SurroundDepth [41]. Resolution refers to the resolution of evaluation. The symbol ${}^{\dagger}$ indicates our implementations. S denotes self-supervised, D denotes depth supervision with Lidar and MS indicates that the scores are averaged per-frame with median-scaling [40, 41] at test time.
Method Resolution Train Test lower is better higher is better
@@
$*{}^{*}$ [60] 1216 $\times$ 1936 S MS 0.362 14.404 14.178 0.412 0.683 0.859 0.922
$*{}^{*}$ [33] 1216 $\times$ 1936 S MS 0.301 5.339 14.115 0.395 0.624 0.828 0.908
Monodepth2 $*{}^{*}$ 1216 $\times$ 1936 S MS 0.217 3.641 12.962 0.323 0.699 0.877 0.939
PackNet-SfM $*{}^{*}$ 1216 $\times$ 1936 S MS 0.234 3.802 13.253 0.331 0.672 0.860 0.931
FSM [40] 1216 $\times$ 1936 S MS 0.202 - - - - - -
$*{}^{*}$ [40] 1216 $\times$ 1936 S MS 0.229 4.589 13.520 0.327 0.677 0.867 0.936
$*{}^{*}$ [41] 1216 $\times$ 1936 S MS 0.200 3.392 12.270 0.301 0.740 0.894 0.947
${}^{\dagger}$ [76] 1216 $\times$ 1936 D - 0.204 3.118 12.330 0.311 0.694 0.872 0.936
${}^{\dagger}$ [35] 1216 $\times$ 1936 D - 0.177 2.525 11.048 0.271 0.747 0.923 0.957
${}^{\dagger}$ [41] 1216 $\times$ 1936 D - 0.175 2.694 11.607 0.283 0.757 0.905 0.953
Ours 1216 $\times$ 1936 D - 0.160 2.527 10.803 0.263 0.799 0.922 0.960
TABLE II: Comparisons for surround-view depth estimation on DDAD dataset [33]. Resolution refers to evaluating resolution. -M indicates occlusion masking. $*{}^{*}$ indicates the implementation of SurroundDepth [41]. The symbol ${}^{\dagger}$ indicates our implementations. S denotes self-supervised, D denotes depth supervision with Lidar and MS indicates that the scores are averaged per-frame with median-scaling [40, 41] at test time.
IV-C Benchmark Results
@@
Results on nuScenes. Because few works study the task of supervised surround-view multi-camera depth estimation, we build a benchmark on the nuScenes [34] dataset. We compare the proposed S3Depth with two methods of self-supervised monocular depth estimation (Monodepth2 [60] and PackNet-SfM [33]), two SOTA self-supervised surround-view depth estimation method (FSM [40] and SurroundDepth [41], as well as two supervised monocular depth estimation (BTS [76] and PixelFormer [35]), as shown in Table I. Specifically, we change the supervision of SurroundDepth [41] from self-supervised to LiDAR supervised and get the result from 0.245 to 0.158 in Abs Rel. It is observed that S3Depth respectively reaches $0.067$ Abs Rel on the evaluation, which exceeds the recently proposed methods: SurroundDepth (Abs Rel $0.158$) and PixelFormer (Abs Rel $0.113$). This demonstrates that S3Depth has the ability to further improve the performance of surround-view depth estimation even while SurroundDepth [41] and PixelFormer [35] have achieved impressive performance.
Figure 4: Qualitative results on DDAD [33] dataset. Top two row are input images and GT depth map. Then, the predicted depth maps are provided. Surr indicates SurroundDepth [41], ${}^{\dagger}$ refers to SurroundDepth [41] supervised with LiDAR. PF denotes PixelFormer [35].
@@
Results on DDAD. We also train and test our model in DDAD [33] datasets, which are known for featuring denser depth estimation in challenging conditions when compared to the nuScenes [34] datasets. As indicated in Table II, our proposed S3Depth also achieves state-of-the-art performance across all evaluation metrics. In addition, we provide qualitative results of different methods on the DDAD [33] dataset in Figure 4. From these results, it is evident that our proposed approach, S3Depth, is capable of predicting more accurate object structures than other competing methods. Our findings suggest that the proposed method can improve the accuracy of depth estimation in challenging conditions and has the potential to be applied in various real-world scenarios.
IV-D Ablation Study
In this subsection, to verify the effects of different modules, we conduct ablation experiments on the nuScenes val set.
Effect of input. The first two rows of the Table III show the effect of different inputs. It yields better results to use surround-view images as input, compared to using random images. This is because surround-view images have overlapping regions which can stabilize gradients, resulting in improved performance. Specifically, the depth estimation results are consistent in the overlap regions between views, which provide additional information that can improve the accuracy of the model’s predictions.
@@ IV-D Ablation Study
Surr. 0.080 0.842 2.882 0.166 0.937 0.961 0.975
Surr. 0.074 0.755 2.708 0.157 0.945 0.964 0.976
@@
Surr. 0.067 0.673 2.457 0.144 0.951 0.970 0.981
TABLE III: The results of depth estimation on nuScenes [34] val set. Adj. stands for Adjacent-view Attention. Random indicates that the input images are selected randomly. Surr. denotes surround-view images captured at the same time
Effect of adjacent-view attention. We conduct the ablation study for adjacent-view attention on the nuScenes [34] datasets. As shown in Table III, the proposed adjacent-attention mechanism facilitates the interaction between adjacent-view information, thereby enabling adjacent-view features to interact with each other. This demonstrates the efficacy of the adjacent-view attention module in enhancing the performance of our model across different backbones. This module effectively enables the network to effectively learn to interact between current and adjacent view features. Moreover, the adjacent-view attention can not only improve the performance but also accelerate the convergence of the model, as shown in Figure 5. We also conduct an ablation study to analyze the impact of the number of adjacent-view attention layers, as shown in Table IV.
Model Z lower is better
@@
MPViT 6 0.074 0.748 2.489 0.147
8 0.067 0.673 2.457 0.144
10 0.071 0.729 2.532 0.149
TABLE IV: The results of depth estimation on nuScenes [34] val set. $Z$ stands for the number of adjacent-view attention layers.
Method lower is better
[…]
L1 loss 0.067 0.632 2.480 0.145
L1+Smooth loss 0.067 0.673 2.457 0.144
Silog loss 0.069 0.502 2.434 0.135
Silog+Smooth loss 0.068 0.485 2.430 0.134
TABLE V: Ablation study of loss on nuScenes [34] val set.
Effect of loss. As shown in Table V, we examine the impact of loss on the performance of our S3Depth. Remarkably, our results indicate that our S3Depth is highly robust to different types of loss functions.
@@
TABLE VI: Ablation study of Feature extraction on Nusense [34] val set. Trans. stands for Transformer. Input size is 640$\times$192, and backbone are pre-trained on ImageNet [79]. $OOM$ refers to Out Of Memory. All backbones are added Adjacent-view Attention module.
Effect of backbone. To delve into the effectiveness of different backbones, we report an ablation study on the nuScenes [34] val set, as shown in Table VI. Also, we report the flops and parameters of different backbones. Benefiting from the module combining CNN and Transformer layers, the MPViT [67] backbone achieves better performance than the pure CNN backbone (ResNet34 [81]). In addition, we further explore the effects of different modules in MPViT [67]. As shown in the bottom of Table VI, both the CNN path and Transformer path play an important role in MPViT [67] for depth estimation.
@@ V Conclusion
In this paper, we build a simple baseline for supervised surround-view depth estimation (S3Depth). Different from monocular depth estimation, the core insight of S3Depth is to exchange adjacent-view information and jointly predict all depths of surrounding views. We use a powerful backbone, which combines CNN with transformer layers for enriched representations. Further, adjacent-view attention is employed at multiple and different scales to incorporate adjacent-view features. Through extensive experiments, we show that S3Depth can achieve state-of-the-art performance on multi-camera depth estimation datasets.

### v4

@@ III-F Loss Function
$L_{depth}=\frac{1}{N_{valid}}\sum_{i}\mathbb{(}{D}^{gt}>0)\left|{D}^{pred}_{i}-{D}^{gt}_{i}\right|,$ (8)
@@
$L_{smooth}=\sum_{x}\left|\partial_{x}d^{\prime}_{x}\right|e^{-\left\|\partial_{x}I_{x}\right\|}+\left|\partial_{y}d^{\prime}_{x}\right|e^{-\left\|\partial_{y}I_{x}\right\|}.$ (9)
@@
$L=L_{depth}+\lambda L_{smooth}.$ (10)
@@ IV Experiments
IV-A Datasets and Evaluation Metrics
NuScenes [10]. The nuScenes benchmark is an up-to-date dataset for 3D object detection, as well as semantic segmentation. It collects 1000 scenes which are divided into 700/150/150 scenes for training/validation/testing. The ground truth depths are generated by processing the LiDAR point cloud. During the training and evaluation procedure, we consider the max distance $D_{max}$ as 80m.
DDAD [9]. The DDAD dataset is a large-scale dataset for long-range (up to 250m) and dense depth estimation in challenging conditions. There are 12,650 training samples (75,900 images) and 3,950 validation samples (15,800 images) in this dataset. During the training and evaluation procedure, we consider the max distance $D_{max}$ as 200m.
Evaluation Metrics. We assess the results of our method following the evaluation protocol of the depth estimation task [2, 9, 20, 21]. While evaluating self-supervised depth estimation methods, the scores are averaged per-frame with median-scaling [20, 21].
@@ IV-B Implementation Details
BTS [13] We use the official implementation of BTS [13]. Densenet161 [70] is used as the backbone. The number of epochs is set to 10 for nuScenes and DDAD.
PixelFormer [11] We use the official implementation of PixelFormer [11]. Swin-Tiny [41] is used as the backbone. The number of epochs is set to 10 for nuScenes and 20 for DDAD.
SurroundDepth [21] We use the official implementation of SurroundDepth [21], and only change the supervision from self-supervised to LiDAR supervised.
@@
TABLE I: Results on the nuScenes[10] val set. indicates the implementation of SurroundDepth [21]. The symbol indicates our implementations. S denotes self-supervised, D denotes depth supervision with LiDAR and MS indicates that the scores are averaged per-frame with median-scaling [20, 21] at test time.
@@
IV-C Comparison to State-of-the-Art
@@
Results on nuScenes. Because few works study the task of supervised surround-view multi-camera depth estimation, we build a benchmark on the nuScenes [10] dataset, as shown in Table I. Specifically, we change the supervision of SurroundDepth [21] from self-supervised to LiDAR supervised and get the result from 0.245 to 0.158 in Abs Rel. AVT-SSDepth respectively reaches $0.067$ Abs Rel, which exceeds the recently proposed methods: SurroundDepth [21] (Abs Rel $0.158$) and PixelFormer [11] (Abs Rel $0.113$). This demonstrates that AVT-SSDepth can further improve the performance of surround-view depth estimation even while SurroundDepth [21] and PixelFormer [11] have achieved impressive performance.
TABLE II: Results on DDAD [9]. -M indicates occlusion masking.
[…]
Monodepth2 [2] S MS 0.362 14.404 14.178 0.683
PackNet-SfM [9] S MS 0.301 5.339 14.115 0.624
Monodepth2 -M S MS 0.217 3.641 12.962 0.699
PackNet-SfM -M S MS 0.234 3.802 13.253 0.672
FSM [20] S MS 0.202 - - -
FSM [20] S MS 0.229 4.589 13.520 0.677
SurroundDepth [21] S MS 0.200 3.392 12.270 0.740
MCDP [73] S MS 0.237 3.030 6.822 0.719
S3D3 [66] S MS 0.169 3.041 11.372 0.809
BTS [13] D - 0.204 3.118 12.330 0.694
PixelFormer [11] D - 0.177 2.525 11.048 0.747
SurroundDepth [21] D - 0.175 2.694 11.607 0.757
Ours D - 0.160 2.527 10.803 0.799
TABLE III: Cross-dataset transfer. Models trained on DDAD are evaluated for scale-aware depth prediction on the NuScenes dataset.
[…]
FSM [20] 0.349 5.064 8.785 0.499
SurroundDepth [21] 0.364 5.476 8.447 0.525
R3D3 [66] 0.292 4.800 7.677 0.660
Ours 0.292 4.201 7.481 0.665
@@
Results on DDAD. We also train and test our model in DDAD [9] datasets, which are known for featuring denser depth estimation in challenging conditions when compared to the nuScenes [10] datasets. As indicated in Table II, our proposed AVT-SSDepth also achieves state-of-the-art performance across all evaluation metrics. From these results, it is evident that our proposed approach, AVT-SSDepth, is capable of predicting more accurate object structures than other competing methods. Our findings suggest that AVT-SSDepth can improve the accuracy of depth estimation in challenging conditions and has the potential to be applied in various real-world scenarios.
Cross-dataset transfer. Compared to previous methods, AVT-SSDepth achieves the lowest Abs Rel (0.292), Sq Rel (4.201), and RMSE (7.481) while attaining the highest $\delta 1$ accuracy (0.665). Notably, it outperforms R3D3 [66], the previous best-performing method, particularly in Sq Rel and RMSE, demonstrating improved generalization across datasets. These results highlight the effectiveness of AVT-SSDepth in adapting to domain shifts, ensuring more reliable depth estimation in diverse driving scenarios.
TABLE IV: Ablation study on nuScenes [10] val set. – Input and Adj. Adj. stands for Adjacent-view Attention. Random indicates that the input images come from diverse time frames. Surr. denotes surround-view images captured at the same time.
@@
ResNet34 [74] Surr. 0.080 0.842 2.882 0.166 0.937 0.961 0.975
ResNet34 [74] Surr. 0.074 0.755 2.708 0.157 0.945 0.964 0.976
@@
MPViT [42] Surr. 0.067 0.673 2.457 0.144 0.951 0.970 0.981
Figure 3: Qualitative results on DDAD [9] dataset. The top two rows are input images and the GT depth map. Then, the predicted depth maps are provided. The last row is RGB point cloud visualization. Surr refers to SurroundDepth [21] supervised with LiDAR. PF denotes PixelFormer [11].
IV-D Ablation Study
Effect of input. The first two rows of the Table IV show the effect of different inputs. It yields better results to use surround-view images as input, compared to using random images. This is because surround-view images have overlapping regions which can stabilize gradients, resulting in improved performance. Specifically, the depth estimation results are consistent in the overlap regions between views, which provide additional information that can improve the accuracy of the model’s predictions. Moreover, this overlapping information can also help the model to better handle changes in perspective, rotation, and scale.
Effect of adjacent-view attention.
As shown in Table IV, the proposed adjacent-attention mechanism facilitates the interaction between adjacent-view information, thereby enabling adjacent-view features to interact with each other. This demonstrates the efficacy of the adjacent-view attention module in enhancing the performance of our model across different backbones. This module effectively enables the network to learn to interact between current and adjacent view features. Moreover, the adjacent-view attention can not only improve the performance but also accelerate the convergence of the model, as shown in Fig.4.
@@ IV-D Ablation Study
TABLE VI: Ablation study on nuScenes [10] val set – $Z$. $Z$ stands for the number of adjacent-view attention layers.
$Z$ lower is better
[…]
4 0.071 0.757 2.507 0.148
6 0.074 0.748 2.489 0.147
8 0.067 0.673 2.457 0.144
10 0.071 0.729 2.532 0.149
TABLE VII: Ablation study of loss on nuScenes [10] val set.
Method lower is better
[…]
L1 loss 0.067 0.632 2.480 0.145
L1+Smooth loss 0.067 0.673 2.457 0.144
Silog loss 0.069 0.502 2.434 0.135
Silog+Smooth loss 0.068 0.485 2.430 0.134
Effect of backbone. To delve into the effectiveness of different backbones, we report an ablation study on the nuScenes [10] val set, as shown in Table V. Also, we report the flops and parameters of different backbones. Benefiting from the module combining CNN and Transformer layers, the MPViT [42] backbone achieves better performance than pure Transformer backbones (PVT-small [75]), the pure CNN backbone (ResNet34 [74]), and hybrid models that combine convolutional networks with transformers (CVT [76]). In addition, we further explore the effects of different modules in MPViT [42]. As shown in the bottom of Table V, both the CNN path and the Transformer path play an important role in MPViT [42] for depth estimation.
Effect of adjacent-view attention layers. We also conduct an ablation study to analyze the impact of the number of adjacent-view attention layers, as shown in Table VI. The model achieves its best results when $Z=8$, highlighting the significance of carefully selecting the number of attention layers
Effect of loss. As shown in Table VII, we examine the impact of loss on the performance of our AVT-SSDepth. Remarkably, our results indicate that our AVT-SSDepth is highly robust to different types of loss functions.
@@ V Conclusion
In this paper, we build a baseline for supervised surround-view depth estimation. Different from monocular depth estimation, the core insight of AVT-SSDepth is to exchange adjacent-view information and jointly predict all depths of surrounding views. We conduct extensive and detailed experiments comparing different backbones. Further, adjacent-view attention is employed at multiple and different scales to incorporate adjacent-view features. Through extensive experiments, we show that AVT-SSDepth can achieve state-of-the-art performance on multi-camera depth estimation datasets. We hope the proposed method can serve as a simple and strong baseline for multi-camera depth estimation.
ACKNOWLEDGMENT
This work was supported by the National Natural Science Foundation of China under Grant 62373356 and the Joint Funds of the National Natural Science Foundation of China under U24B20162.